• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Mazouz

    Log entry time 00:04:20 on October 7,2004

    Entry number 133228

    keyword=Shift summary

    Shift summary

    19:40 Beam is back. HARP Scan OK.
    20:00 Compton studies (Empty target)
    21:10 Ask MCC for 2.5 uA. (LH2 Target). Turn off the Proton Array HV for the elastic runs.
    21:45 beam is back after 30min of RF problems. Bleedthrough check is 2.05%
    22:30 Start of run 3224 (DVCSNoProton configuration). ARS Saturation in the calo.
    11:30 Start of run 3227.